2. Calculate the AH for the reaction:




CS2 (1) + 2 O2 (g) CO2 (g) + 2 SO2 (g)




Given:




AHCO2 (g) = -393.5 kJ/




mol;




AHSO2 = -296.8 kJ/mol;




AHCS2 (1) = 87.9kJ/mol

1
Expert's answer
2022-03-14T18:18:03-0400

The given reaction is "CS_2(l)+ 2O_2(g) \longrightarrow CO_2(g)+2SO_2(g)"

"\Delta H_f (CO_2)=-393.5 \frac{kJ}{mol}"

"\Delta H_f(SO_2)=-296.8 kJ/mol"

"\Delta H_f( CS_2)=87.9 kJ/mol"

"\Delta H_r= \Delta H_f(CO_2)+\Delta H_f(SO_2)-\Delta H_f(CS_2)"



"\Delta H_r= -393.5-296.8-87.9"

"\Delta H_r= -778.2 \frac{kJ}{mol}"